Father's age is six times his son's age. Nine years hence the age of father will be three times his son's age. Find their present ages.
Answer by stanbon(75887)

Father's age is six times his son's age. Nine years hence the age of father will be three times his son's age. Find their present ages.
-----
Equations:
f = 6s
f+9 = 3(s+9)
-----
Substitute for "f" and solve for "s":
-----
6s+9 = 3s + 27
----
3s = 18
----
s = 6
=============
Cheers,
Stan H.
=============